Consider two Stackelberg firms, Firm Alpha and Firm Omega, each with marginal costs of 50 and each facing the market inverse demand curve: P=500-1/2Q Firm Alpha moves first, Firm Omega moves second. How many MORE units does Firm Alpha produce than Firm Omega due to first mover advantage?
Answer #1

DAILY PLAN Alpha (leaders) Omega (follower) profit Total Total Revenue Cost Jo = 1500-6462) Q2 50 Q2 2 Q2 Output produced byDATE 2 Ja=5006 - 6 © (450-Q, (45o- 2 Solo 2 2 2 dura dQ 500-20 450 +22 2 ។ -So=o 2 225 @ + ☺ 2 225= Qi @ 2 Tu5o= 6 Output pro